### v2

# A Uniform Rewriting Algorithm for Twisted Tensor Representations of Finite General Linear Groups

Abstract.
Let $q=p^{f}$ be a prime power and let $H$ be a classical group of type $A$, so that $H\cong\mathrm{SL}_{d}(q)$ or $\mathrm{GL}_{d}(q)$, acting on its natural module $V$ of dimension $d$ over $\mathbb{F}_{q}$. Suppose $W$ is an absolutely irreducible $\mathbb{F}_{q}H$–module such that, over an algebraic closure, it is isomorphic to a twisted tensor product
$W\;\cong\;\bigotimes_{t=1}^{r}L(\lambda^{(t)})^{(e_{t})},$
where each $L(\lambda^{(t)})$ is an irreducible polynomial representation of $\mathrm{GL}_{d}$ of degree $k_{t}$, the integers $e_{t}\geq 0$ denote Frobenius twists, and the total polynomial degree $K=\sum_{t=1}^{r}k_{t}$ satisfies K<q1K<q-1. Since Frobenius twists merely apply field automorphisms to eigenvalues and do not merge distinct eigenvalues or change eigenspace dimensions, the spectral analysis can be carried out on the underlying untwisted tensor product $\bigotimes_{t=1}^{r}L(\lambda^{(t)})$.
Although constructive recognition of classical groups is well developed, there is currently no general, uniform rewriting algorithm for such twisted tensor products: existing handlers either treat only very small highest weights, or work in low overall dimension, and rely on ad hoc eigenvalue calculations. To address this limitation, we prove a general base-$q$ injectivity lemma which implies that, for a Singer cycle $s\in H$, the eigenvalues of $s$ on $W\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{d}}$ are parametrised by digit vectors of length $d$ with entries bounded by $K$. In particular, distinct weights of $W$ correspond to distinct eigenvalues. If in addition $W$ is multiplicity-free for the diagonal torus of $\mathrm{GL}_{d}$, then every eigenspace of $s$ is one-dimensional, so $s$ has a simple spectrum on $W$. This gives a uniform conceptual explanation of the eigenvalue computations in work of Gül–Ankaralıoğlu, which was previously restricted to a fixed list of small highest weights. Using this spectral property, we design a Las Vegas algorithm which, given $G\leq\mathrm{GL}(W)$ isomorphic to $H$ and acting irreducibly on $W$, constructs a projective representation of $G$ of degree $d$ equivalent to the natural representation of $H$. The expected running time is polynomial in $d$, $\log q$, $K$, the cost of field operations in $\mathbb{F}_{q^{d}}$, and the cost of discrete logarithm computations in $\mathbb{F}_{q^{d}}^{\times}$. Our algorithm extends the rewriting algorithms of Magaard–O’Brien–Seress and of Gül–Ankaralıoğlu to a wider class of twisted tensor products of polynomial modules. We also discuss practical aspects and include SageMath code illustrating the core subroutines.
Key words and phrases:
Matrix group recognition, Rewriting algorithm, Twisted tensor product, Polynomial representation, Singer cycle
2020 Mathematics Subject Classification:
20C40, 20G05, 20G40
1. Introduction
1.1. Constructive recognition of matrix groups
Constructive recognition of finite groups is a central theme in computational group theory. Given a finite group $G$ specified in some implicit form, for example as a subgroup of a permutation group or a matrix group, one aims to construct an explicit isomorphism between $G$ and a standard copy of a known abstract group $H$. This problem has been intensively studied for symmetric and alternating groups, classical groups, and other families of groups of Lie type; see, for example, the survey of Beals–Leedham-Green–Niemeyer–Praeger–Seress for symmetric and alternating groups, and Brooksbank’s work on classical groups in their natural representation [1].
In the setting of matrix groups, the natural representation plays a distinguished role. For a classical group $H$ of dimension $d$ over $\mathbb{F}_{q}$, a great deal of structure is visible in its action on the natural module $V\cong\mathbb{F}_{q}^{d}$. Explicit recognition algorithms for $H\leq\mathrm{GL}(V)$, in the natural representation, were developed by Brooksbank and others [1], and later extended to the black–box setting and more general classical groups by Dietrich–Leedham-Green–O’Brien [3]. These algorithms typically assume that the given group acts on a space of dimension $d$ and that the representation is already natural (or close to natural).
In practice, however, matrix groups often arise via representations of dimension $n$ different from $d$, and a central task is to rewrite such a representation in terms of the natural one. Formally, suppose $G\leq\mathrm{GL}(W)$ is a group generated by a set of matrices $X$ acting irreducibly on an nn–dimensional $\mathbb{F}_{q}$–vector space $W$, and that $G$ is known to be isomorphic to a classical group $H$ of rank $d$. The rewriting problem is to construct a homomorphism
$\varphi:G\longrightarrow\mathrm{GL}(V),$
where $V$ is the natural dd–dimensional module for $H$, such that $\varphi$ is an isomorphism onto a subgroup between $\mathrm{SL}_{d}(q)$ and $\mathrm{GL}_{d}(q)$, and such that the map can be computed algorithmically on group elements.
1.2. Previous work on rewriting algorithms
The first general treatment of rewriting for small-dimensional representations is due to Magaard, O’Brien and Seress [7]. They consider the case when $G\cong H$ with $\mathrm{SL}_{d}(q)\leq H\leq\mathrm{GL}_{d}(q)$ and $W$ is an irreducible $\mathbb{F}_{q}G$–module of dimension at most $d^{2}$. They provide a Las Vegas polynomial-time algorithm which constructs a projective representation of degree $d$ for $G$ from its action on $W$. The algorithm relies heavily on the spectral properties of certain elements and a detailed analysis of the possible modules $W$ of dimension at most $d^{2}$.
Subsequently, more specialised rewriting algorithms have been developed for particular classes of representations that occur frequently in computational practice. Corr [2] treats the symmetric square representation: he presents and analyses a Las Vegas algorithm for rewriting the representation afforded by $\mathrm{Sym}^{2}(V)$ to a projective copy of the natural representation. This yields performance improvements in situations where such symmetric square representations appear as composition factors.
A further step was taken by Gül and Ankaralıoğlu [4]. They study the case where $W$ is in a tensor family of twisted modules of degree between $d^{2}$ and $d^{3}$. More precisely, they assume $W$ is a twisted tensor product of highest weight modules with highest weights among
$\lambda_{1},\lambda_{2},\lambda_{d-2},\lambda_{d-1},2\lambda_{1},2\lambda_{d-1}$
for a classical group $H$ of type $A$, and they develop a Las Vegas algorithm that rewrites the action on $W$ to a projective action of degree $d$. Their analysis combines representation theory (via Steinberg’s tensor product theorem) with careful, yet ad-hoc, eigenvalue computations for particular tensor products such as
$V\otimes V^{\tau}\otimes V^{\tau^{2}},\quad V\otimes(\wedge^{2}V)^{\tau},\quad V\otimes(\mathrm{Sym}^{2}V)^{\tau},$
where $\tau$ is a Frobenius automorphism.
These contributions fit into the broader matrix group recognition project, which aims to build a general framework for the constructive recognition of finite matrix groups via composition trees and local handlers for particular types of composition factors and modules; see, for example, [3, 1] and references therein.
1.3. Limitations of existing work
The rewriting algorithms in [7, 4, 2] are precise and effective for their intended families of modules, but their scope is restricted in two ways.
First, the work of Magaard–O’Brien–Seress is constrained to representations of dimension at most $d^{2}$. Beyond this range, the classification of possible irreducible modules becomes significantly more complex. Their methods rely on a detailed understanding of the small-degree representation theory of $\mathrm{GL}_{d}$ in defining characteristic and do not immediately extend to larger families of modules.
Second, the twisted-module algorithm of Gül–Ankaralıoğlu [4] focuses on a fixed list of highest weights of small polynomial degree. The proofs involve a case-by-case analysis of the eigenvalues of Singer cycles on each of the corresponding tensor products. While this approach works well for the particular modules considered, it does not directly generalise to arbitrary polynomial highest weights or more complicated tensor constructions.
On the other hand, the general black–box recognition algorithms for classical groups [3, 1] treat all representations uniformly, without exploiting representation-theoretic structure such as the polynomial degree of highest weights. This leads to algorithms that are broadly applicable but may be suboptimal on specific families of modules.
1.4. Contribution of this paper
The aim of this paper is to address these limitations by introducing a structural condition under which Singer cycles have a simple spectrum on tensor products of polynomial modules, and by exploiting this to obtain a general rewriting algorithm.
We focus on groups of type $A$, that is, $H\cong\mathrm{SL}_{d}(q)$ or $\mathrm{GL}_{d}(q)$, and consider irreducible modules $W$ which, over an algebraic closure, can be written as a twisted tensor product
$W\;\cong\;\bigotimes_{t=1}^{r}L(\lambda^{(t)})^{(e_{t})},$
where $L(\lambda^{(t)})$ is the irreducible highest weight module with highest weight $\lambda^{(t)}$ and the integers $e_{t}\geq 0$ denote Frobenius twists. Since Frobenius twists by powers of $q$ merely apply field automorphisms to eigenvalues and do not merge distinct eigenvalues or change eigenspace dimensions (see Section 2), our spectral analysis may (and will) be carried out on the underlying untwisted tensor product
$\bigotimes_{t=1}^{r}L(\lambda^{(t)}),$
suppressing the twists from the notation.
Assuming that the $L(\lambda^{(t)})$ are polynomial representations of degrees $k_{t}$ (in the standard sense for $\mathrm{GL}_{d}$) and that their total degree
$K:=\sum_{t=1}^{r}k_{t}$
satisfies K<q1K<q-1, we first prove a general base-$q$ injectivity lemma which implies:
  • for a fixed Singer cycle $s$, distinct weights of $W$ correspond to distinct eigenvalues of $s$ on $W\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{d}}$;
  • if in addition $W$ is multiplicity-free for the diagonal torus, then all eigenspaces of $s$ are one-dimensional (simple spectrum).
The key input is a simple number-theoretic lemma: if one writes the exponent of an eigenvalue as an integer in base $q$, then under the constraint K<q1K<q-1 the digits (which encode weight multiplicities) are uniquely determined modulo $q^{d}-1$. This yields a uniform generalisation of the eigenvalue computations in [4], avoiding case-by-case analysis of particular highest weights.
Building on this spectral property, we design a general rewriting algorithm for representations in this class. Our main constructive result is formulated as follows.
Suppose $G\leq\mathrm{GL}(W)$ is a group of matrices over $\mathbb{F}_{q}$, generated by a set $X$, such that $G$ acts irreducibly on $W$ and is isomorphic to a group $H$ satisfying $\mathrm{SL}_{d}(q)\leq H\leq\mathrm{GL}_{d}(q)$. Assume that (over an algebraic closure) $W$ is isomorphic to a twisted tensor product of polynomial representations of total degree K<q1K<q-1 which is multiplicity-free for the diagonal torus. Then there exists a Las Vegas algorithm which, given $X$, constructs a projective representation of degree $d$ of $G$ that is equivalent to the natural representation of $H$. The running time is polynomial in $d$, $\log q$, the cost of field operations in $\mathbb{F}_{q^{d}}$, the cost of generating nearly uniform random elements of $G$, and the cost of discrete logarithm computations in $\mathbb{F}_{q^{d}}^{\times}$.
Within the framework of the Matrix Group Recognition Project, this algorithm serves as a specialized handler that significantly expands the class of amenable representations. By replacing the ad-hoc eigenvalue analysis of previous works with a uniform number-theoretic approach, we strictly extend the scope of the algorithm of Gül–Ankaralıoğlu [4] to arbitrary polynomial highest weights (subject to K<q1K<q-1), while maintaining a polynomial time complexity comparable to the specific cases previously known. For a detailed contextualization of this result within the existing literature and a discussion of specific limitations, we refer the reader to Remark 5.2 in Section 5.
The paper is organised as follows. In Section 2 we collect notation and recall basic facts about polynomial representations of $\mathrm{GL}_{d}(q)$ and tensor products. Section 3 contains the number-theoretic injectivity lemma which underlies our spectral analysis. Section 4 establishes the distinct eigenvalue property and the simple spectrum property under multiplicity-freeness. Section 5 describes the rewriting algorithm and proves its correctness and complexity bounds. Section 6 presents illustrative SageMath code for core subroutines. Finally, Section 7 reports on computational experiments that verify the base-$q$ injectivity lemma and demonstrate the simple spectrum property for symmetric powers of the natural module.
2. Preliminaries
2.1. Finite fields and classical groups of type A
Throughout, $p$ denotes a prime and $q=p^{f}$ a prime power, with $f\geq 1$. We write $\mathbb{F}_{q}$ for the finite field of order $q$ and $\mathbb{F}_{q^{d}}$ for its extension of degree $d$. The multiplicative group of $\mathbb{F}_{q^{d}}$ is cyclic of order $q^{d}-1$.
Let $d\geq 2$ and let $V$ be a dd–dimensional vector space over $\mathbb{F}_{q}$. We write $\mathrm{GL}_{d}(q)$ for the group of invertible linear transformations of $V$, and $\mathrm{SL}_{d}(q)$ for the subgroup of determinant 1 transformations. We fix a basis of $V$ and identify $\mathrm{GL}_{d}(q)$ with the group of invertible $d\times d$ matrices over $\mathbb{F}_{q}$. We also write $\mathrm{PGL}_{d}(q)$ for the projective general linear group $\mathrm{GL}_{d}(q)/Z(\mathrm{GL}_{d}(q))$.
Definition 2.1.
A Singer cycle in $\mathrm{GL}_{d}(q)$ is an element whose eigenvalues on $V\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{d}}$ form a single orbit under the qq–Frobenius automorphism; equivalently, an element conjugate to the companion matrix of an irreducible polynomial of degree $d$ over $\mathbb{F}_{q}$.
Concretely, let $\omega$ be a generator of $\mathbb{F}_{q^{d}}^{\times}$. Then there exists a basis of $V\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{d}}$ with respect to which a Singer cycle $s$ acts diagonally as
$s\cdot e_{i}=\ell_{i}e_{i},\quad\ell_{i}=\omega^{q^{i-1}},\qquad i=1,\dots,d.$
We shall assume such a Singer cycle is contained in $H$, where $H$ denotes a group satisfying $\mathrm{SL}_{d}(q)\leq H\leq\mathrm{GL}_{d}(q)$.
2.2. Polynomial representations and highest weights
We recall basic facts about polynomial representations of $\mathrm{GL}_{d}$ over fields of positive characteristic. Our main reference is Jantzen’s monograph: see [5, Part II].
Let $K$ be an algebraic closure of $\mathbb{F}_{q}$, and view $\mathrm{GL}_{d}=\mathrm{GL}(V_{K})$ as an algebraic group over $K$, where $V_{K}=V\otimes_{\mathbb{F}_{q}}K$. Thus, in this subsection $\mathrm{GL}_{d}$ denotes the algebraic group over $K$, while $\mathrm{GL}_{d}(q)$ from the previous subsection is its group of $\mathbb{F}_{q}$–rational points. A rational representation of $\mathrm{GL}_{d}$ is called polynomial of degree $k$ if it is realised as a subrepresentation of the kk–fold tensor power $V_{K}^{\otimes k}$ and the action is given by substitution of matrices in homogeneous polynomials of degree $k$ in the matrix entries.
Irreducible rational representations of $\mathrm{GL}_{d}$ are parametrised by dominant weights $\lambda=(\lambda_{1},\dots,\lambda_{d})$ with integers $\lambda_{1}\geq\cdots\geq\lambda_{d}$. When all $\lambda_{i}$ are non-negative, the corresponding module $L(\lambda)$ is polynomial and has degree $k(\lambda):=\lambda_{1}+\cdots+\lambda_{d}$; see, e.g., [5]. For the restriction to the algebraic subgroup $\mathrm{SL}_{d}\subset\mathrm{GL}_{d}$, we may regard $\lambda$ modulo multiples of $(1,1,\dots,1)$, but this plays no role in our arguments.
Definition 2.2.
Let $\lambda$ be a dominant weight with non-negative entries. The polynomial degree of $L(\lambda)$ is
$k(\lambda)=\sum_{i=1}^{d}\lambda_{i}.$
For a finite list $\lambda^{(1)},\dots,\lambda^{(r)}$ of such weights, we set
$K:=\sum_{t=1}^{r}k(\lambda^{(t)}).$
We shall need a simple description of the weights of $L(\lambda)$ when $\lambda$ is polynomial.
Let $T$ denote the diagonal torus of $\mathrm{GL}_{d}$, consisting of elements of the form $\mathrm{diag}(t_{1},\dots,t_{d})$ with $t_{i}\in K^{\times}$. Let $\varepsilon_{i}:T\to K^{\times}$ be the character given by $\varepsilon_{i}(\mathrm{diag}(t_{1},\dots,t_{d}))=t_{i}$. Every weight $\mu$ of a rational $\mathrm{GL}_{d}$–module can be expressed as
$\mu=\sum_{i=1}^{d}b_{i}(\mu)\varepsilon_{i}$
with $b_{i}(\mu)\in\mathbb{Z}$.
Proposition 2.3.
Let $L(\lambda)$ be an irreducible polynomial $\mathrm{GL}_{d}$–module of degree $k=k(\lambda)$, and let $\mu$ be a weight of $L(\lambda)$ with respect to $T$. Then
$b_{i}(\mu)\in\mathbb{Z}_{\geq 0}\quad\text{for all }i,\qquad\sum_{i=1}^{d}b_{i}(\mu)=k.$
Proof.
This is standard; we only sketch the idea and refer to [5] for details. Polynomial representations of $\mathrm{GL}_{d}$ correspond to representations of the Schur algebra and can be realised as direct summands of $V_{K}^{\otimes k}$. In $V_{K}^{\otimes k}$ a pure tensor $e_{i_{1}}\otimes\cdots\otimes e_{i_{k}}$ has weight $\varepsilon_{i_{1}}+\cdots+\varepsilon_{i_{k}}$, so each weight has the form $\sum b_{i}\varepsilon_{i}$ with $b_{i}\geq 0$ and $\sum b_{i}=k$. Weights of any polynomial representation then arise as those of the ambient tensor power, with the same form. ∎
In particular, we may regard each weight $\mu$ of $L(\lambda)$ as encoded by a vector $b(\mu)=(b_{1}(\mu),\dots,b_{d}(\mu))$ in
$\mathcal{B}_{k}:=\Bigl\{(b_{1},\dots,b_{d})\in\mathbb{Z}_{\geq 0}^{d}\;\Bigm|\;\sum_{i=1}^{d}b_{i}=k\Bigr\}.$
Definition 2.4 (Multiplicity-free polynomial module).
Let $L(\lambda)$ be a rational polynomial $\mathrm{GL}_{d}$–module. We say that $L(\lambda)$ is multiplicity-free (for the diagonal torus $T$) if every weight of $L(\lambda)$ with respect to $T$ occurs with multiplicity $1$, i.e. every weight space is one-dimensional. More generally, a finite tensor product $W=\bigotimes_{t=1}^{r}L(\lambda^{(t)})$ is called multiplicity-free if each of its weights (as a TT–module) occurs with multiplicity $1$.
Important examples include the symmetric powers $\mathrm{Sym}^{k}(V)$ and exterior powers $\wedge^{k}V$ of the natural module $V$; in these cases each weight is determined uniquely by the multiset of indices and hence occurs with multiplicity $1$.
2.3. Frobenius twists and tensor products
Let $G=\mathrm{GL}_{d}$ considered as an algebraic group over $\overline{\mathbb{F}}_{q}$. Let $F:G\to G$ denote the Frobenius morphism defining $G$ over $\mathbb{F}_{q}$. On diagonal torus elements $t=\mathrm{diag}(t_{1},\dots,t_{d})$ we have
$F(t)=\mathrm{diag}(t_{1}^{\,q},\dots,t_{d}^{\,q}).$
Given a rational KGKG–module $M$, the FF–twist $M^{(1)}$ is defined by composing the action of $G$ on $M$ with $F$; more generally, for an integer $r\geq 0$ we define the rr–th Frobenius twist $M^{(r)}$ by composing with $F^{r}$. On weight spaces, this has the effect of raising eigenvalues to $q^{r}$–th powers.
In this paper the spectral analysis is carried out for untwisted tensor products $W=\bigotimes_{t}L(\lambda^{(t)})$. Frobenius twists by powers of $q$ merely apply field automorphisms to eigenvalues and permute them within their Frobenius orbits; they do not merge distinct eigenvalues or change the dimensions of eigenspaces. Thus, for the purpose of eigenvalue distinctness and eigenspace dimensions, it suffices to consider untwisted tensor products. We therefore omit twists in the main spectral arguments and discuss them only at the level of module structure.
Steinberg’s tensor product theorem describes irreducible rational representations in terms of Frobenius twists and pp–restricted weights; see [5, Part II, §3]. We only need it as structural background and will not use it explicitly in proofs.
2.4. The rewriting problem for polynomial tensor modules
Let $H$ be a finite group satisfying $\mathrm{SL}_{d}(q)\leq H\leq\mathrm{GL}_{d}(q)$, acting on its natural module $V$ over $\mathbb{F}_{q}$. We are given a subgroup $G\leq\mathrm{GL}(W)$, with $W$ an nn–dimensional vector space over $\mathbb{F}_{q}$, such that:
  • $G$ acts irreducibly on $W$;
  • $G$ is isomorphic to $H$ (via some unknown isomorphism);
  • Over an algebraic closure $\overline{\mathbb{F}}_{q}$, the $\overline{\mathbb{F}}_{q}G$–module $W\otimes_{\mathbb{F}_{q}}\overline{\mathbb{F}}_{q}$ is isomorphic to a tensor product
    $\bigotimes_{t=1}^{r}L(\lambda^{(t)}),$
    where each $L(\lambda^{(t)})$ is an irreducible polynomial representation of $\mathrm{GL}_{d}$ of degree $k_{t}$, and $K:=\sum_{t=1}^{r}k_{t}<q-1$.
We make the standard algorithmic assumptions that:
  • we have access to the generating set $X$ of $G$ as matrices in $\mathrm{GL}_{n}(\mathbb{F}_{q})$;
  • we can multiply matrices and compute in $\mathbb{F}_{q}$ (and in $\mathbb{F}_{q^{d}}$ when needed);
  • we can sample nearly uniform random elements of $G$ at cost $\xi$ per element.
The rewriting problem in this context is:
Problem. Construct a projective representation
$\varphi:G\longrightarrow\mathrm{PGL}_{d}(q)$
that is equivalent to the natural projective representation of $H$ on $V$, and such that $\varphi(g)$ can be effectively computed from the matrix of $g$ on $W$.
Our goal is to design a Las Vegas algorithm solving this problem, under the additional hypothesis that $W$ is multiplicity-free (Definition 2.4), with running time polynomial in $d$, $\log q$, $K$, and the cost of matrix and field operations.
3. A number-theoretic injectivity lemma
The key technical ingredient in our analysis is the following elementary lemma about writing integers in base $q$. It asserts that, under a suitable bound on the digits, the map from digit vectors to residues modulo $q^{d}-1$ is injective.
Lemma 3.1 (Base-$q$ injectivity).
Let $q\geq 2$, $d\geq 1$ and let $C$ be an integer with $0\leq C<q-1$. Consider the set
$\mathcal{B}_{C}\;=\;\bigl\{\mathbf{b}=(b_{1},\dots,b_{d})\in\mathbb{Z}_{\geq 0}^{d}\bigm|0\leq b_{i}\leq C\text{ for all }i\bigr\}$
and the map
$\Phi:\mathcal{B}_{C}\longrightarrow\mathbb{Z}/(q^{d}-1)\mathbb{Z},\qquad\Phi(\mathbf{b})=\sum_{i=1}^{d}b_{i}q^{i-1}\bmod(q^{d}-1).$
Then $\Phi$ is injective.
Proof.
For $\mathbf{b}\in\mathcal{B}_{C}$ we have
$0\leq\Phi(\mathbf{b})\leq C\sum_{i=1}^{d}q^{i-1}=C\,\frac{q^{d}-1}{q-1}.$
Since C<q1C<q-1, it follows that
$C\,\frac{q^{d}-1}{q-1}<(q-1)\,\frac{q^{d}-1}{q-1}=q^{d}-1.$
Thus $\Phi(\mathbf{b})$ is actually an integer in $0,q^{d}-2]$.
Suppose $\mathbf{b},\mathbf{c}\in\mathcal{B}_{C}$ with $\Phi(\mathbf{b})\equiv\Phi(\mathbf{c})\pmod{q^{d}-1}$. Then there exists $k\in\mathbb{Z}$ such that
$\Phi(\mathbf{b})-\Phi(\mathbf{c})=k(q^{d}-1).$
The left-hand side lies in the interval $(-(q^{d}-1),q^{d}-1)$, so necessarily $k=0$. Hence $\Phi(\mathbf{b})=\Phi(\mathbf{c})$ as integers.
Now view $\Phi(\mathbf{b})$ and $\Phi(\mathbf{c})$ as base-$q$ expansions:
$\Phi(\mathbf{b})=b_{1}+b_{2}q+\cdots+b_{d}q^{d-1},\qquad\Phi(\mathbf{c})=c_{1}+c_{2}q+\cdots+c_{d}q^{d-1}.$
Since $0\leq b_{i},c_{i}\leq C\leq q-2<q$, each expression is the base-$q$ representation of an integer with digits in $\{0,\dots,q-1\}$. The base-$q$ representation is unique, so $b_{i}=c_{i}$ for all $i$, and hence $\mathbf{b}=\mathbf{c}$. Thus $\Phi$ is injective. ∎
This lemma will be applied to the weight multiplicities of polynomial modules, which will play the role of the digits $b_{i}$.
4. Distinct eigenvalues for tensor products
We now combine Lemma 3.1 with the weight structure of polynomial modules to establish a general distinct-eigenvalue property for Singer cycles on tensor products, and a simple spectrum result under multiplicity-freeness.
4.1. Eigenvalues of a Singer cycle on a polynomial module
Let $s$ be a Singer cycle in $H$. Over $\mathbb{F}_{q^{d}}$ we may choose a basis of $V\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{d}}$ such that $s$ acts as
$s\cdot e_{i}=\ell_{i}e_{i},\qquad\ell_{i}=\omega^{q^{i-1}},\quad i=1,\dots,d,$
for some generator $\omega$ of $\mathbb{F}_{q^{d}}^{\times}$.
Let $L(\lambda)$ be an irreducible polynomial representation of $\mathrm{GL}_{d}$ of degree $k=k(\lambda)$, realised over $K$. Let $T$ be the diagonal torus as above, and let $\mu$ be a weight of $L(\lambda)$ with weight vector $b(\mu)=(b_{1}(\mu),\dots,b_{d}(\mu))$ as in Proposition 2.3. Then for $t=\mathrm{diag}(t_{1},\dots,t_{d})\in T$ the action on a weight vector of weight $\mu$ is
$t\cdot v=\biggl(\prod_{i=1}^{d}t_{i}^{\,b_{i}(\mu)}\biggr)v.$
In particular, if we regard $s$ as an element of $T$ over $K$, then $s$ acts on this weight space with eigenvalue
$\prod_{i=1}^{d}\ell_{i}^{\,b_{i}(\mu)}=\prod_{i=1}^{d}\omega^{b_{i}(\mu)q^{i-1}}=\omega^{E(\mu)},$
where
(4.1) $E(\mu)=\sum_{i=1}^{d}b_{i}(\mu)q^{i-1}\in\mathbb{Z}/(q^{d}-1)\mathbb{Z}.$
By Proposition 2.3 we have $b_{i}(\mu)\geq 0$ and $\sum_{i}b_{i}(\mu)=k$, so $b_{i}(\mu)\in[0,k]$.
4.2. Eigenvalues on tensor products
Let $\lambda^{(1)},\dots,\lambda^{(r)}$ be dominant weights with non-negative entries, and suppose $L(\lambda^{(t)})$ is polynomial of degree $k_{t}=k(\lambda^{(t)})$. Consider the tensor product
$W=\bigotimes_{t=1}^{r}L(\lambda^{(t)}).$
As a module for the diagonal torus $T$, the weights of $W$ are sums
$\nu=\mu^{(1)}+\cdots+\mu^{(r)},$
where $\mu^{(t)}$ is a weight of $L(\lambda^{(t)})$. Let $b^{(t)}(\mu^{(t)})=(b_{1}^{(t)},\dots,b_{d}^{(t)})$ denote the corresponding weight vector, so that
$\mu^{(t)}=\sum_{i=1}^{d}b_{i}^{(t)}\varepsilon_{i}.$
Definition 4.1.
For a weight $\nu$ of $W$ as above, define
$c_{i}(\nu)=\sum_{t=1}^{r}b_{i}^{(t)},\qquad\mathbf{c}(\nu)=(c_{1}(\nu),\dots,c_{d}(\nu)).$
Then $c_{i}(\nu)\geq 0$ and
$\sum_{i=1}^{d}c_{i}(\nu)=\sum_{t=1}^{r}\sum_{i=1}^{d}b_{i}^{(t)}=\sum_{t=1}^{r}k_{t}=K.$
Thus $\mathbf{c}(\nu)\in\mathcal{B}_{K}$, where $K$ is the total polynomial degree.
The eigenvalue of $s$ on the weight space of $\mu^{(t)}$ in $L(\lambda^{(t)})$ is $\omega^{E(\mu^{(t)})}$ with
$E(\mu^{(t)})=\sum_{i=1}^{d}b_{i}^{(t)}q^{i-1}.$
Therefore, the eigenvalue of $s$ on a pure tensor
$v_{1}\otimes\cdots\otimes v_{r}\in W$
with $v_{t}$ in the weight space of $\mu^{(t)}$ is
$\omega^{E(\mu^{(1)})}\cdots\omega^{E(\mu^{(r)})}=\omega^{E(\nu)},\qquad E(\nu)=\sum_{t=1}^{r}E(\mu^{(t)})=\sum_{i=1}^{d}c_{i}(\nu)q^{i-1}.$
Hence the eigenvalue of $s$ on the weight space of $\nu$ in $W$ is $\omega^{E(\nu)}$, where $E(\nu)$ depends only on $\mathbf{c}(\nu)$.
4.3. Distinct eigenvalues for distinct weights
We can now state and prove the first main spectral result.
Theorem 4.2 (Distinct eigenvalues for different weights).
Let $H$ be a classical group of type $A$ with $\mathrm{SL}_{d}(q)\leq H\leq\mathrm{GL}_{d}(q)$, and let $s$ be a Singer cycle in $H$. Let $\lambda^{(1)},\dots,\lambda^{(r)}$ be dominant weights with non-negative entries such that each $L(\lambda^{(t)})$ is polynomial of degree $k_{t}$, and set $K=\sum_{t=1}^{r}k_{t}$. Assume K<q1K<q-1.
Let
$W=\bigotimes_{t=1}^{r}L(\lambda^{(t)})$
be the corresponding tensor product. Then for any two distinct weights $\nu\neq\nu^{\prime}$ of $W$ (viewed as characters of $T$), the eigenvalues of $s$ on the weight spaces of $\nu$ and $\nu^{\prime}$ are distinct.
Proof.
Let $\nu$ be a weight of $W$, with associated vector $\mathbf{c}(\nu)\in\mathcal{B}_{K}$ and exponent
$E(\nu)=\sum_{i=1}^{d}c_{i}(\nu)q^{i-1}\in\mathbb{Z}/(q^{d}-1)\mathbb{Z}.$
If $\nu\neq\nu^{\prime}$ then, as weights, the corresponding vectors $c_{i}(\nu)$ and $c_{i}(\nu^{\prime})$ differ for some $i$. Thus $\mathbf{c}(\nu)\neq\mathbf{c}(\nu^{\prime})$ as elements of $\mathcal{B}_{K}$.
By Lemma 3.1 with $C=K$, the map
$\mathbf{c}\mapsto\sum_{i=1}^{d}c_{i}q^{i-1}\pmod{q^{d}-1}$
is injective on $\mathcal{B}_{K}$. Hence $E(\nu)\neq E(\nu^{\prime})$ in $\mathbb{Z}/(q^{d}-1)\mathbb{Z}$. It follows that the eigenvalues of $s$ on the corresponding weight spaces are distinct in $\mathbb{F}_{q^{d}}^{\times}$. ∎
Remark 4.3.
In all applications in this paper, we are interested in modules of positive total polynomial degree $K>0$. The hypothesis K<q1K<q-1 then forces $q\geq 3$. Indeed, if $q=2$ then K<q1K<q-1 implies $K<1$ and hence $K=0$, so there are no non-trivial polynomial tensor products satisfying our standing assumption. Thus Theorem 4.2 and Theorem 4.4 below are non-vacuous only for $q\geq 3$. Moreover, Theorem 4.2 is a statement about distinct weights (characters) of $T$; it does not address the dimensions of the corresponding weight spaces, which may be larger than one in general.
4.4. Simple spectrum under multiplicity-freeness
We now state the strengthened result under the additional assumption that $W$ is multiplicity-free.
Theorem 4.4 (Simple spectrum under multiplicity-freeness).
In the setting of Theorem 4.2, assume in addition that the tensor product
$W=\bigotimes_{t=1}^{r}L(\lambda^{(t)})$
is multiplicity-free for the diagonal torus $T$ (Definition 2.4). Then every eigenspace of $s$ on $W\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{d}}$ is one-dimensional. Equivalently, $s$ has a simple spectrum on $W$.
Proof.
As $s$ lies in the diagonal torus (after extension of scalars), each weight space $W_{\nu}$ of $W$ is contained in an eigenspace of $s$. By multiplicity-freeness, each weight space $W_{\nu}$ is one-dimensional. Moreover, Theorem 4.2 tells us that distinct weights give distinct eigenvalues. Thus each eigenspace of $s$ coincides with a unique one-dimensional weight space. ∎
Remark 4.5.
The modules considered by Gül and Ankaralıoğlu in [4] are twisted tensor products of modules with highest weights among
$\lambda_{1},\lambda_{2},\lambda_{d-2},\lambda_{d-1},2\lambda_{1},2\lambda_{d-1}.$
These all have polynomial degree at most $2$, and the corresponding untwisted modules are multiplicity-free for $T$. In their setting, at most three such factors are tensored, so $K\leq 6$. For $q\geq 8$, the condition K<q1K<q-1 is automatically satisfied, and Theorem 4.4 recovers their simple-spectrum property in a uniform way. In small fields, some additional low-degree exceptions must be treated separately, as is done explicitly in [4].
5. A general rewriting algorithm
In this section we describe a rewriting algorithm for representations in the class covered by Theorem 4.4. We sketch the main steps and then formalise the result as a Las Vegas algorithm with complexity bounds.
5.1. Outline of the algorithm
Let $G\leq\mathrm{GL}(W)$ be as in the problem statement of Section 2, with $W$ an irreducible $\mathbb{F}_{q}G$–module and $W\otimes\overline{\mathbb{F}}_{q}$ isomorphic to a tensor product of multiplicity-free polynomial modules of total degree K<q1K<q-1.
Step 1: Find a Singer-type element.
Using standard nearly uniform random element generators (e.g. product replacement), we sample elements $s\in G$ until we find one with the following properties:
  • the minimal polynomial of $s$ on $W$ has a large irreducible factor of degree $d$;
  • the order of $s$ is divisible by a primitive prime divisor of $q^{d}-1$.
The existing analysis in [7, 4] shows that the proportion of such elements is bounded below by a positive constant independent of $d$ and $q$, apart from a finite list of small exceptions. Thus, after $O(\log\varepsilon^{-1})$ trials we find such an element with probability at least $1-\varepsilon$. By Theorem 4.4, this $s$ has a simple spectrum on $W$.
We factor the characteristic (or minimal) polynomial of $s$ over $\mathbb{F}_{q}$ and, using standard algorithms, determine a root $\omega$ in $\mathbb{F}_{q^{d}}$ corresponding to the eigenvalues of $s$ of degree $d$. From this, we obtain all eigenvalues of $s$ on $W$ as elements of $\mathbb{F}_{q^{d}}$.
Step 2: Label eigenvalues via base-$q$ expansions.
Let $\lambda_{1},\dots,\lambda_{n}$ be the eigenvalues of $s$ on $W$, written as $\lambda_{j}=\omega^{E_{j}}$ with $E_{j}\in\{0,1,\dots,q^{d}-2\}$. Write $E_{j}$ in base $q$:
$E_{j}=c_{1}^{(j)}+c_{2}^{(j)}q+\cdots+c_{d}^{(j)}q^{d-1}.$
By Proposition 2.3 and Theorem 4.2, in our situation the digits actually satisfy $0\leq c_{i}^{(j)}\leq K$ and $\sum_{i}c_{i}^{(j)}=K.$ By Lemma 3.1 and the discussion in Section 4, each eigenvalue corresponds to a unique vector
$\mathbf{c}^{(j)}=\bigl(c_{1}^{(j)},\dots,c_{d}^{(j)}\bigr)\in\mathcal{B}_{K}.$
These vectors $\mathbf{c}^{(j)}$ represent the combined multiplicities of the basic eigenvalues $\ell_{i}$ on each tensor factor.
Step 3: Choose a tensor-compatible eigenbasis of $W$.
For each eigenvalue $\lambda_{j}$ we compute an eigenvector $f_{j}\in W\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{d}}$. Since each eigenspace is one-dimensional by Theorem 4.4, a scalar multiple of $f_{j}$ is uniquely determined up to multiplication by $\mathbb{F}_{q^{d}}^{\times}$. We normalise each $f_{j}$ in a consistent way (for example, fixing a coordinate equal to 1) and regard $\{f_{1},\dots,f_{n}\}$ as a basis of $W\otimes\mathbb{F}_{q^{d}}$.
The association $j\mapsto\mathbf{c}^{(j)}$ then provides a labelling of this basis by elements of $\mathcal{B}_{K}$. Moreover, by comparing with the combinatorics of the weight sets of the factors $L(\lambda^{(t)})$ (for example, via semistandard tableaux), we can identify which $\mathbf{c}^{(j)}$ correspond to basis vectors that project to a fixed basis of $V$.
In particular, when the tensor factors are symmetric or exterior powers of $V$, the labels $\mathbf{c}^{(j)}$ directly encode multiplicities of basis vectors of $V$ and allow us to reconstruct a basis of $V$ inside $W\otimes\mathbb{F}_{q^{d}}$.
Step 4: Recover the action of $G$ on $V$.
Let $g\in G$ and let $M_{W}(g)$ be the matrix of $g$ on $W$ with respect to the eigenbasis $\{f_{j}\}$. Given that $W$ is a tensor product of polynomial modules, the matrix entries of $M_{W}(g)$ are polynomial expressions in the entries of the matrix $A(g)$ of $g$ on $V$. For example, if $W=V\otimes V$, then
$M_{W}(g)=A(g)\otimes A(g),$
so each entry of $M_{W}(g)$ is a product $a_{i\ell}\,a_{jm}$ in terms of entries of $A(g)$.
More generally, for each factor $L(\lambda^{(t)})$, the matrix of $g$ on $L(\lambda^{(t)})$ is given by a known polynomial functor (Schur functor) applied to $A(g)$. The matrix $M_{W}(g)$ on $W$ is then the tensor product of these matrices.
Our strategy is to choose suitable entries of $M_{W}(g)$ that correspond to monomials with few variables and to solve for the entries of $A(g)$ from these equations. This is analogous to Lemma 4.1 of [4] and to the analysis in [2] for the symmetric square.
Concretely, we proceed as follows:
  • Identify a subset of basis vectors $\{f_{j_{1}},\dots,f_{j_{m}}\}$ whose labels $\mathbf{c}^{(j)}$ correspond to weight patterns that make the corresponding matrix entries in $M_{W}(g)$ particularly simple (e.g. single monomials in a small set of $a_{uv}$).
  • For each such pair $(j_{u},j_{v})$ we compute the coefficient $k_{uv}$ of $f_{j_{v}}$ in the expansion of $g\cdot f_{j_{u}}$.
  • Each equation $k_{uv}$ yields a relation among the entries of $A(g)$.
  • By choosing enough such equations, we obtain a system from which we can solve for $A(g)$ up to scalar multiplication and field automorphisms.
The details depend on the specific list of highest weights $\lambda^{(t)}$, but the general principle is that the monomial structure of the polynomial functor describing $W$ allows us to isolate individual entries of $A(g)$.
Step 5: Assemble a projective representation.
Once we can construct the matrix $A(g)$ for each $g\in X$, we obtain a map
$\varphi:G\longrightarrow\mathrm{GL}_{d}(\mathbb{F}_{q^{d}})$
by extending multiplicatively. The image of $\varphi$ is contained in a group between $\mathrm{SL}_{d}(q)$ and $\mathrm{GL}_{d}(q)$. Standard techniques then allow us to descend $\varphi$ to a projective representation over $\mathbb{F}_{q}$ and to normalise it so that it agrees with a standard copy of $H$.
5.2. Main rewriting theorem
We now state the main algorithmic result formally. Throughout this section we treat arithmetic in $\mathbb{F}_{q^{d}}$, the generation of nearly uniform random elements of $G$, and discrete logarithm computations in $\mathbb{F}_{q^{d}}^{\times}$ as basic subroutines. All complexity bounds are measured in terms of the cost of field operations in $\mathbb{F}_{q^{d}}$, the cost $\xi$ of generating random elements of $G$, and the cost $\delta_{q^{d}}$ of discrete logarithm computations in $\mathbb{F}_{q^{d}}^{\times}$.
Theorem 5.1 (Rewriting algorithm for multiplicity-free polynomial tensor modules).
Let $q=p^{f}$ be a prime power and let $H$ be a group with $\mathrm{SL}_{d}(q)\leq H\leq\mathrm{GL}_{d}(q)$. Let $V$ be the natural dd–dimensional $\mathbb{F}_{q}H$–module. Suppose $G\leq\mathrm{GL}(W)$ is a group of matrices over $\mathbb{F}_{q}$, acting irreducibly on the nn–dimensional $\mathbb{F}_{q}$–vector space $W$, and that $G\cong H$.
Assume that over $\overline{\mathbb{F}}_{q}$ the module $W\otimes_{\mathbb{F}_{q}}\overline{\mathbb{F}}_{q}$ is isomorphic to a tensor product
$\bigotimes_{t=1}^{r}L(\lambda^{(t)}),$
where each $L(\lambda^{(t)})$ is an irreducible polynomial representation of $\mathrm{GL}_{d}$ of degree $k_{t}$, $K=\sum_{t=1}^{r}k_{t}<q-1$, and the tensor product is multiplicity-free for the diagonal torus (Definition 2.4). We also assume that the polynomial functors defining each factor $L(\lambda^{(t)})$ are known explicitly (for example, each $L(\lambda^{(t)})$ is specified as a Schur functor such as a symmetric or exterior power of $V$).
Assume further that we can:
  • sample nearly uniform random elements of $G$ at cost $\xi$ per element;
  • perform arithmetic operations in $\mathbb{F}_{q^{d}}$ at cost $\rho_{q^{d}}$ per operation;
  • compute discrete logarithms in $\mathbb{F}_{q^{d}}^{\times}$ at cost $\delta_{q^{d}}$ per instance.
Then there exists a Las Vegas algorithm which, given a generating set $X$ for $G$, with probability at least $1-\varepsilon$:
  • constructs a projective representation $\varphi:G\to\mathrm{PGL}_{d}(q)$ equivalent to the natural projective representation of $H$ on $V$;
  • outputs explicit matrices $\varphi(x)$ for all $x\in X$.
The expected running time of the algorithm is polynomial in $d$, $\log q$, $K$, $\log(\varepsilon^{-1})$, and in $\xi$, $\rho_{q^{d}}$ and $\delta_{q^{d}}$.
Proof.
The algorithm is as outlined above. We briefly indicate how the pieces fit together and why the running time is polynomial, referring to the analysis in [7, 4, 2] for standard subroutines.
(1) Finding a Singer-type element. The methods of [7, 4] show that a random element of $G$ has a primitive prime divisor of $q^{d}-1$ in its order with probability bounded below by a positive constant, apart from finitely many small pairs $(d,q)$. For classical groups of type $A$, this can be achieved by considering elements whose characteristic polynomial has an irreducible factor of degree $d$. Factoring characteristic polynomials and testing degrees can be done in time polynomial in $d$ and $\log q$. After $O(\log\varepsilon^{-1})$ samples, we find such an element $s$ with probability at least $1-\varepsilon$. By Theorem 4.4, $s$ has a simple spectrum on $W$.
(2) Eigenvalues and base-$q$ labelling. We compute the characteristic polynomial of $s$ on $W$ and factor it over $\mathbb{F}_{q}$. An irreducible factor of degree $d$ yields an extension field isomorphic to $\mathbb{F}_{q^{d}}$, in which we can express a chosen root as $\omega$. The eigenvalues of $s$ on $W$ are then obtained by factoring the minimal polynomial of $s$ and computing roots in $\mathbb{F}_{q^{d}}$. This step is polynomial in $n$, $d$ and $\log q$.
We write each eigenvalue $\lambda_{j}$ as $\omega^{E_{j}}$ with $0\leq E_{j}\leq q^{d}-2$ and expand $E_{j}$ in base $q$:
$E_{j}=c_{1}^{(j)}+c_{2}^{(j)}q+\cdots+c_{d}^{(j)}q^{d-1}.$
By Lemma 3.1, this yields a unique vector $\mathbf{c}^{(j)}\in\mathcal{B}_{K}$.
(3) Constructing an eigenbasis. For each eigenvalue $\lambda_{j}$ we compute an eigenvector $f_{j}$ by standard linear algebra over $\mathbb{F}_{q^{d}}$. Since each eigenspace is one-dimensional (Theorem 4.4), this requires solving a homogeneous system of linear equations of size $n$ over $\mathbb{F}_{q^{d}}$, at cost $O(n^{3})$ field operations. Thus we obtain an eigenbasis $\{f_{j}\}$ of $W\otimes\mathbb{F}_{q^{d}}$ together with labels $\mathbf{c}^{(j)}$.
(4) Recovering the action on $V$. Given a generator $x\in X$, we compute the matrix $M_{W}(x)$ of $x$ with respect to the eigenbasis $\{f_{j}\}$ by expressing $xf_{j}$ as a linear combination of the $f_{k}$’s. Each application of $x$ to a basis vector and subsequent decomposition costs $O(n^{2})$ operations in $\mathbb{F}_{q^{d}}$, so constructing $M_{W}(x)$ costs $O(n^{3})$ field operations.
We then select a set of matrix entries $k_{uv}$ of $M_{W}(x)$ for which the corresponding monomials in the entries of $A(x)$ are simple and can be identified via the combinatorics of the weight labels $\mathbf{c}^{(j)}$. The precise selection depends on the highest weights $\lambda^{(t)}$, but in all cases we can choose $O(d^{2})$ entries to obtain a system of polynomial equations in the $d^{2}$ unknown entries of $A(x)$.
These equations can be solved using algebraic techniques similar to those in [4, 2], often by taking ratios to eliminate scalars and using the structure of the Schur functor to isolate individual $a_{ij}$. Since the degrees of the polynomials are bounded in terms of $K$, and $K$ is fixed (or at least small compared to $q$), this system can be solved in time polynomial in $d$ and $K$.
The result is a matrix $A(x)\in\mathrm{GL}_{d}(\mathbb{F}_{q^{d}})$ determined up to a scalar and possibly up to a field automorphism. Consistent choices across generators yield a representation
$\varphi:G\longrightarrow\mathrm{GL}_{d}(\mathbb{F}_{q^{d}}).$
(5) Projectivisation and descent. The map $\varphi$ is a homomorphism because the matrices $M_{W}(g)$ respect the group structure and the extraction of $A(g)$ is functorial in $g$. Composing with the natural projection $\mathrm{GL}_{d}\to\mathrm{PGL}_{d}$ gives a projective representation. By a standard argument (e.g. [1, 3]) one can adjust $\varphi$ by scalars and field automorphisms to ensure that its image lies in $\mathrm{GL}_{d}(q)$ and coincides with a standard copy of $H$.
Each of these steps has expected running time bounded by a polynomial in $d,\log q,K,$ and in the costs $\xi$ and $\rho_{q^{d}}$. The dependence on $\varepsilon$ is logarithmic via the number of trials needed in Step 1. Thus the overall algorithm is Las Vegas with the stated complexity. ∎
Remark 5.2 (Comparison with existing algorithms and limitations).
From the perspective of matrix group recognition, the algorithm of Theorem 5.1 functions as a dedicated module-specific handler. It is designed to integrate seamlessly with and enhance current generic recognition algorithms rather than replacing them.
First, Theorem 5.1 strictly extends the scope of the twisted-module algorithm of Gül–Ankaralıoğlu [4], which treats only tensor products built from a fixed list of small highest weights. Our condition K<q1K<q-1 allows arbitrary polynomial highest weights subject to a mild total degree bound, and the eigenvalue analysis follows uniformly from the general base-$q$ injectivity lemma rather than from case-by-case manipulation of particular tensor products. In particular, specialised handlers such as Corr’s algorithm for the symmetric square representation [2] arise as low-degree instances (with $K=2$) of our general polynomial-tensor setting. The asymptotic running time remains similar to [4], being dominated by the computation of eigenvalues and eigenvectors for a Singer cycle.
Second, there is a useful comparison with the small-dimensional handler of Magaard–O’Brien–Seress [7]. Their algorithm works for arbitrary irreducible $\mathbb{F}_{q}G$–modules of dimension at most $d^{2}$, without assuming any polynomial highest-weight structure, but relies on a detailed classification of such modules and does not directly extend beyond this range. In contrast, our method requires strong representation-theoretic structure (a twisted tensor product of polynomial modules of total degree K<q1K<q-1 which is multiplicity-free for the diagonal torus), but imposes no a priori upper bound on $\dim W$; in particular, it applies to many tensor products whose dimension is significantly larger than $d^{2}$. Thus the two approaches are complementary: [7] handles small-dimensional modules with no structural assumptions, while Theorem 5.1 provides a uniform treatment of a broad class of polynomial tensor modules of bounded degree.
Third, the constructive recognition algorithms of Brooksbank [1] and of Dietrich–Leedham-Green–O’Brien [3] address a different layer of the recognition pipeline. They assume that the group already acts on (or close to) its natural module and then construct an explicit isomorphism with a standard copy of the corresponding classical group, often in a black-box setting. Once our rewriting step has been carried out and a projective copy of the natural representation has been recovered, these algorithms can be applied directly to obtain a full constructive isomorphism between $G$ and a standard copy of $H$. In this sense, Theorem 5.1 extends the library of available handlers that feed into the existing natural-module and black-box recognition frameworks.
On the other hand, the condition K<q1K<q-1 excludes certain very small fields (for instance, $q\leq 4$ when $K=3$) where the base-$q$ injectivity property fails. Previous works such as [4] treat these low-field configurations via explicit case-by-case analysis, whereas our approach prioritises a uniform algorithm for generic $q$. We note that the excluded cases where $K\geq q-1$ typically involve very small fields. In such restricted settings, alternative strategies such as brute-force search or precomputed lookup tables are computationally feasible and effective; thus, this condition does not constitute a significant barrier in practice. Additionally, if the module $W$ is not multiplicity-free, a Singer cycle does not completely separate the eigenspaces, and the simple-spectrum argument breaks down. In such cases it is natural to combine our eigenvalue-based labelling with MeatAxe or algebra-based methods, as in [7], to further decompose the resulting high-dimensional eigenspaces.
6. Illustrative SageMath code
In this section we present some SageMath code fragments that illustrate core components of the algorithm: computing eigenvalues of a Singer cycle, extracting base-$q$ expansions, and checking the injectivity property of Lemma 3.1 for a given module.
6.1. Base-$q$ expansion and injectivity test
The following function takes an exponent $E$ and returns its base-$q$ expansion in $d$ digits. We assume $0\leq E<q^{d}$.
1def base_q_expansion(E, q, d):
2 """
3 Return the base-q expansion of integer E as a list of length d:
[…]
5 """
6 coeffs = []
7 for _ in range(d):
8 coeffs.append(E % q)
9 E //= q
10 return coeffs # c[0], ..., c[d-1]
We can use this to verify the injectivity of the map $\Phi$ for given parameters $(q,d,C)$:
1def check_injectivity(q, d, C, verbose=True):
2 r"""
3 Check injectivity of the map
4
5 Phi : B_C -> Z/(q^d - 1)Z,
[…]
7
8 where
[…]
10
11 This is an exhaustive search, so exponential in d.
12 Intended only for small d and C.
13 """
14 from itertools import product
15
16 modulus = q**d - 1
17 seen = {}
18
19 for b in product(range(C + 1), repeat=d):
20 E = sum(b[i] * (q**i) for i in range(d)) % modulus
21 if E in seen and seen[E] != b:
22 if verbose:
23 print("Collision found!")
24 print(" b =", b, "and", "c =", seen[E], "both map to", E)
25 return False
26 seen[E] = b
27
28 if verbose:
[…]
30 print(f"Checked {len(seen)} vectors.")
31 return True
For small values of $d$ and $C$ one can experimentally confirm Lemma 3.1.
Note. The implementation above uses itertools.product which is suitable only for small parameters. For the large-scale experiments in Section 7.2 below (e.g., $q=2^{16},d=10$), our supplementary script employs an optimized recursive generator weight_patterns_sumK and performs arithmetic purely on integer exponents modulo $q^{d}-1$, avoiding the expensive construction of the extension field.
6.2. Eigenvalues of a Singer cycle on a tensor power
As a simple model, we consider the action on $V^{\otimes K}$. In practice, $W$ is a submodule or quotient of $V^{\otimes K}$ corresponding to a Schur functor, but $V^{\otimes K}$ is easier to work with.
1def singer_eigenvalues_tensor_power(q, d, K):
2 """
[…]
4 where V is the natural d-dimensional module over GF(q).
5
6 Returns a dictionary mapping base-q exponent vectors c in B_K
7 (with sum(c) = K) to the corresponding eigenvalue in GF(q^d).
8
9 Note: This function verifies the injectivity of the map
[…]
12 It does NOT detect possible weight multiplicities in irreducible
13 submodules or quotients: in such modules several linearly
14 independent vectors may share the same weight pattern c and
15 hence the same eigenvalue.
16 """
17 # Finite fields
18 Fq = GF(q)
19 Fqd = GF(q**d, ’a’)
20 a = Fqd.gen()
21
[…]
23 omega = a # usually a is primitive
24
25 # Dictionary from c-vector to eigenvalue
26 eig = {}
27
28 # Enumerate all c in B_K
29 from itertools import product
30
31 for c in product(range(K+1), repeat=d):
32 if sum(c) != K:
33 continue
34 # exponent E(c) = sum c_i q^i
35 E = sum(c[i] * (q**i) for i in range(d))
36 eig_val = omega**E
37 eig[c] = eig_val
38
39 return eig
By inspecting the dictionary returned by singer_eigenvalues_tensor_power, one can verify that different $\mathbf{c}$ give distinct eigenvalues when K<q1K<q-1, in agreement with Lemma 3.1.
6.3. Recovering base-$q$ labels from eigenvalues
Suppose we know an eigenvalue $\lambda\in\mathbb{F}_{q^{d}}$ and we have fixed a generator $\omega$ of $\mathbb{F}_{q^{d}}^{\times}$. We can compute the exponent $E$ such that $\lambda=\omega^{E}$, and then extract its base-$q$ expansion. The following function performs this task.
1def exponent_and_digits(lam, omega, q, d):
2 """
3 Given an eigenvalue lam = omega^E in GF(q^d)^*,
4 return the exponent E (0 <= E < q^d-1) and its base-q digits.
5 """
6 # Discrete log: find E such that omega^E = lam
7 E = lam.log(omega) # Sage’s discrete log
8 E = ZZ(E) % (q**d - 1) # normalise exponent modulo q^d - 1
9 digits = base_q_expansion(E, q, d)
10 return E, digits
This function provides the labels $\mathbf{c}^{(j)}$ used in Step 2 of the algorithm.
Besides these basic routines, the full script singer_sym_check.sage (archived with this paper) also contains helper functions for constructing an explicit Singer matrix in $\mathrm{GL}_{d}(q)$, computing the induced action on $\mathrm{Sym}^{k}(V)$, running end-to-end eigenvalue checks, and performing the toy reconstruction experiment from $\mathrm{Sym}^{2}(A)$ described in Section 7 below. For readability, we omit these longer code fragments here.
7. Computational experiments
In this section, we report on some small computational experiments which serve as a sanity check for the number-theoretic Lemma 3.1 and for the spectral behaviour of Singer cycles on symmetric powers of the natural module, as well as a toy model for the reconstruction step in the rewriting algorithm. All computations were performed in SageMath; the corresponding script is available at https://github.com/phucdv2018/singer_sym_check.sage (archived at DOI: https://doi.org/10.5281/zenodo.17792033) and included as a supplementary file singer_sym_check.sage. The code is primarily intended as a collection of sanity checks for the key mechanisms in the paper. While the matrix reconstruction routines are designed for small examples to illustrate correctness, the eigenvalue labeling and injectivity checks are implemented efficiently and scale to large parameters.
7.1. Verification of the base-$q$ injectivity lemma
Recall that Lemma 3.1 asserts that, for C<q1C<q-1, the map
$\Phi:\mathcal{B}_{C}\to\mathbb{Z}/(q^{d}-1)\mathbb{Z},\qquad\Phi(\mathbf{b})=\sum_{i=1}^{d}b_{i}q^{i-1}\bmod(q^{d}-1),$
is injective, where
$\mathcal{B}_{C}=\{(b_{1},\dots,b_{d})\in\mathbb{Z}_{\geq 0}^{d}\mid 0\leq b_{i}\leq C\text{ for all }i\}.$
For fixed parameters $(q,d,C)$, the script exhaustively enumerates $\mathbf{b}\in\mathcal{B}_{C}$, computes $\Phi(\mathbf{b})$ and checks for collisions. Since this is exponential in $d$, we only run it for small values. For example, with
$q=7,\quad d=3,\quad C=3,$
we have C<q1C<q-1 and $|\mathcal{B}_{3}|=4^{3}=64$. The program confirms that
$\Phi:\mathcal{B}_{3}\longrightarrow\mathbb{Z}/(7^{3}-1)\mathbb{Z}$
is injective on all $64$ vectors and reports no collisions. In particular, a typical run outputs
Phi is injective on B_3 for (q,d,C) = (7,3,3).
followed by
Checked 64 vectors.
Similar calculations for various small triples $(q,d,C)$ with C<q1C<q-1 consistently support Lemma 3.1. Of course, these experiments do not constitute a proof, but they provide additional confidence that the base-$q$ injectivity phenomenon behaves as predicted in all small cases.
7.2. Model eigenvalues on tensor powers
Before turning to genuine matrix representations, we also implemented the abstract eigenvalue model described in Section 4. Given parameters $(q,d,K)$ with K<q1K<q-1, we consider
$\mathcal{B}_{K}^{\prime}=\{\mathbf{c}=(c_{1},\dots,c_{d})\in\mathbb{Z}_{\geq 0}^{d}\mid\sum_{i}c_{i}=K\},$
and define, for each $\mathbf{c}\in\mathcal{B}_{K}^{\prime}$,
$E(\mathbf{c})=\sum_{i=1}^{d}c_{i}q^{i-1},\qquad\lambda_{\mathbf{c}}=\omega^{E(\mathbf{c})}\in\mathbb{F}_{q^{d}}^{\times},$
where $\omega$ is a fixed generator of $\mathbb{F}_{q^{d}}^{\times}$. This matches the expression for the eigenvalues of a Singer cycle on $V^{\otimes K}$ (and on polynomial submodules) given in (4.1).
For $(q,d,K)=(7,3,3)$ the set $\mathcal{B}_{3}^{\prime}$ has size $|\mathcal{B}_{3}^{\prime}|=\binom{3+3-1}{3-1}=10$, corresponding to the $10$ ways of writing $3$ as an ordered sum of $3$ non-negative integers. The script computes $\lambda_{\mathbf{c}}$ for all $\mathbf{c}\in\mathcal{B}_{3}^{\prime}$ and verifies that these $10$ eigenvalues are pairwise distinct. This is reported in the console as
Number of weight patterns c with sum(c) = 3: 10
followed by
Distinct weight patterns c give distinct eigenvalues (as expected).
Moreover, for a sample pattern, say $\mathbf{c}=(0,0,3)$, the program obtains an eigenvalue
$\lambda_{\mathbf{c}}=\omega^{147},\qquad 147=0+0\cdot 7+3\cdot 7^{2},$
and recovers the base-$7$ digits of $147$ as $(0,0,3)$. In the actual output this appears as
Example weight pattern c = (0, 0, 3),
Exponent E = 147,    Base-q digits = [0, 0, 3].
This exactly illustrates the mechanism of Lemma 3.1: the bound K<q1K<q-1 ensures that the digits $c_{i}$ can be reconstructed uniquely from the exponent $E(\mathbf{c})$ modulo $q^{d}-1$.
As a more demanding sanity check closer to the intended applications, we also ran a purely “exponent–model” variant of this experiment for larger parameters. Specifically, we considered
$(q,d,K)\;=\;(2^{16},\,10,\,K)\qquad\text{with}\qquad K\in\{2,3,4\}.$
In this setting we do not construct $\mathbb{F}_{q^{d}}$ or a Singer matrix explicitly, but work only with the exponents $E(\mathbf{c})$ in $\mathbb{Z}/(q^{d}-1)\mathbb{Z}$. For each $K$ we enumerate all digit vectors $\mathbf{c}\in\mathcal{B}_{K}$ of length $d$ with $\sum_{i}c_{i}=K$ and verify that the map
$\mathbf{c}\;\longmapsto\;\sum_{i=1}^{d}c_{i}q^{i-1}\bmod(q^{d}-1)$
is injective. For $K=2,3,4$ we have
$|\mathcal{B}_{K}^{\prime}|\;=\;\binom{d+K-1}{K}\;=\;55,\ 220,\ 715,\ \ \mbox{respectively},$
and in each case the program reports that all $|\mathcal{B}_{K}^{\prime}|$ exponents are distinct and that no collisions occur. This confirms the base-$q$ injectivity lemma in a regime where the module dimension reaches
$\dim\mathrm{Sym}^{4}(V)\;=\;\binom{10+4-1}{4}\;=\;715,$
while avoiding any explicit matrix computations over the large field $\mathbb{F}_{2^{160}}$.
Finally, we demonstrate the feasibility of our labeling strategy for these parameters. Our implementation generates the lookup table for $K=4$ and verifies injectivity modulo $2^{160}-1$ in approximately $30$ milliseconds on a standard workstation. This confirms that, by working directly with exponents and look-up tables rather than discrete logarithms, the eigenvalue identification step has negligible computational cost compared to matrix operations, even over very large fields.
7.3. Real Singer matrices and symmetric powers
To test the full representation-theoretic picture, we next worked with genuine matrices in $\mathrm{GL}_{d}(q)$ and their induced action on symmetric powers of the natural module.
Fix $(q,d)=(7,3)$ and let $\mathbb{F}_{7^{3}}$ be the field of order $7^{3}$. Let $\omega$ be a generator of $\mathbb{F}_{7^{3}}^{\times}$ and consider $\mathbb{F}_{7^{3}}$ as a 33–dimensional vector space over $\mathbb{F}_{7}$ with basis $\{1,\omega,\omega^{2}\}$. Multiplication by $\omega$ defines a linear operator $m_{\omega}$ on $\mathbb{F}_{7^{3}}$, and with respect to this basis we obtain a matrix $S\in\mathrm{GL}_{3}(7)$ which is a Singer cycle. For one particular choice of irreducible polynomial over $\mathbb{F}_{7}$, the script produces the companion matrix
$S=\begin{pmatrix}0&0&3\\ 1&0&0\\ 0&1&1\end{pmatrix}$,
which is therefore a Singer cycle in $\mathrm{GL}_{3}(7)$ by Definition 2.1; this matrix also appears explicitly in the console output.
We then construct the induced matrices of $S$ on $\mathrm{Sym}^{k}(V)$, where $V\cong\mathbb{F}_{7}^{3}$ is the natural module and $k=2,3$. This is done concretely via the tensor power $V^{\otimes k}$ and the symmetrisation operator: we work with the basis of $V^{\otimes k}$ indexed by kk–tuples of $\{0,1,2\}$, apply the kk–fold tensor product $S^{\otimes k}$, and restrict to the symmetric subspace spanned by symmetrised basis vectors. The resulting matrices $S^{(k)}\in\mathrm{GL}(\mathrm{Sym}^{k}(V))$ have dimensions
$\dim\mathrm{Sym}^{2}(V)=\binom{3+2-1}{2}=6,\qquad\dim\mathrm{Sym}^{3}(V)=\binom{3+3-1}{3}=10,$
as expected, and the script prints these dimensions together with an explicit list of the multi-indices labelling the basis of $\mathrm{Sym}^{k}(V)$.
Over $\mathbb{F}_{7^{3}}$, we compute the eigenvalues of $S^{(k)}$ and compare them with the theoretical eigenvalues
$\lambda_{\mathbf{c}}=\omega^{\sum_{i}c_{i}7^{i-1}},\qquad\mathbf{c}=(c_{1},c_{2},c_{3})\in\mathbb{Z}_{\geq 0}^{3},\ \sum_{i}c_{i}=k.$
The experiments yield the following:
  • For $k=2$: there are exactly $6$ distinct eigenvalues of $S^{(2)}$ on $\mathrm{Sym}^{2}(V)$, and they coincide with the $6$ values $\lambda_{\mathbf{c}}$ arising from the $6$ patterns $\mathbf{c}$ with $\sum c_{i}=2$.
  • For $k=3$: there are exactly $10$ distinct eigenvalues of $S^{(3)}$ on $\mathrm{Sym}^{3}(V)$, and they coincide with the $10$ values $\lambda_{\mathbf{c}}$ arising from the $10$ patterns $\mathbf{c}$ with $\sum c_{i}=3$.
In both cases the script also checks the multiplicities of eigenvalues. For $k=2$ it reports
Number of eigenvalues returned (with multiplicity): 6
followed by
All eigenvalues have multiplicity 1 (simple spectrum).
Similarly, for $k=3$ it reports
Number of eigenvalues returned (with multiplicity): 10
and again
All eigenvalues have multiplicity 1 (simple spectrum).
Thus, in these examples $S^{(2)}$ and $S^{(3)}$ both have simple spectrum on $\mathrm{Sym}^{2}(V)$ and $\mathrm{Sym}^{3}(V)$, respectively. This is consistent with the fact that these modules are multiplicity-free for the diagonal torus, and provides concrete examples illustrating Theorem 4.4 in the simplest non-trivial cases.
Finally, the script compares the sets of eigenvalues obtained from the actual matrices $S^{(k)}$ with the model set $\{\lambda_{\mathbf{c}}:\mathbf{c}\in\mathcal{B}_{k}^{\prime},\,\sum_{i}c_{i}=k\}$. In both cases it prints
$\texttt{Size of set of eigenvalues (real) :}\quad\dim\mathrm{Sym}^{k}(V),$
$\texttt{Size of set of eigenvalues (model) :}\quad|\mathcal{B}_{k}^{\prime}|,$
followed by
SUCCESS: Eigenvalues on Sym^k(V) match the digit-vector model.
Since $|\mathcal{B}_{k}^{\prime}|=\dim\mathrm{Sym}^{k}(V)$, this confirms that the distinct eigenvalues are in bijection with the weight patterns $\mathbf{c}$, exactly as predicted by Theorem 4.2.
As a typical example in the case $k=3$, the script reports an eigenvalue $\lambda=\omega^{147}$ for $S^{(3)}$. Taking discrete logarithms and expanding $147$ in base $7$ yields digits $(0,0,3)$. This demonstrates that the eigenvalue indeed corresponds to the weight pattern $\mathbf{c}=(0,0,3)$, in perfect agreement with the formula in Section 4.
7.4. A toy reconstruction experiment for the rewriting step
The rewriting algorithm of Section 5 relies, in Step 4, on recovering the underlying matrix $A(g)$ of an element $g$ on the natural module $V$ from its action on a polynomial module such as a symmetric or exterior power. In full generality, this is done by identifying suitable matrix entries of $M_{W}(g)$ (on $W$) which are given by low-degree monomials in the entries of $A(g)$ and solving the resulting system of polynomial equations. To illustrate this mechanism in a particularly simple setting, we implemented a toy model for the symmetric square in dimension $d=2$.
Let $V=\mathbb{F}_{q}^{2}$ with $q$ odd, and let $e_{0},e_{1}$ be the standard basis of $V$. We identify $\mathrm{Sym}^{2}(V)$ with the 33–dimensional space spanned by
$v_{1}=e_{0}\otimes e_{0},\quad v_{2}=e_{0}\otimes e_{1}+e_{1}\otimes e_{0},\quad v_{3}=e_{1}\otimes e_{1}.$
For a matrix
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{GL}_{2}(q),$
we let $A$ act on $V^{\otimes 2}$ diagonally via
$A\cdot(u\otimes w)=(Au)\otimes(Aw).$
In particular,
$Ae_{0}=ae_{0}+ce_{1},\qquad Ae_{1}=be_{0}+de_{1}.$
We briefly spell out the computation for $v_{1}$; the other cases are analogous. Using linearity of the tensor product,
$A\cdot v_{1}\;=\;A\cdot(e_{0}\otimes e_{0})\;=\;(Ae_{0})\otimes(Ae_{0})\;=\;(ae_{0}+ce_{1})\otimes(ae_{0}+ce_{1}),$
so
$(ae_{0}+ce_{1})\otimes(ae_{0}+ce_{1})=a^{2}(e_{0}\otimes e_{0})+ac(e_{0}\otimes e_{1}+e_{1}\otimes e_{0})+c^{2}(e_{1}\otimes e_{1})=a^{2}v_{1}+ac\,v_{2}+c^{2}v_{3}.$
A similar expansion for $v_{2}$ and $v_{3}$ yields
$A\cdot v_{3}=b^{2}v_{1}+bd\,v_{2}+d^{2}v_{3},$
$A\cdot v_{2}=2ab\,v_{1}+2(ad+bc)\,v_{2}+2cd\,v_{3}.$
Thus, with respect to the basis $\{v_{1},v_{2},v_{3}\}$, the matrix $\mathrm{Sym}^{2}(A)$ has columns
$\mathrm{Sym}^{2}(A)=\begin{pmatrix}a^{2}&2ab&b^{2}\\ ac&2(ad+bc)&bd\\ c^{2}&2cd&d^{2}\end{pmatrix}$.
The script contains a function that implements this formula and, given a matrix $M\in\mathrm{GL}_{3}(q)$ of this form, performs a brute-force search over all $A\in\mathrm{GL}_{2}(q)$ to find those satisfying $\mathrm{Sym}^{2}(A)=M$. Since the field is finite and we restrict ourselves to $q=7$, this is perfectly feasible for experimental purposes. The aim is to verify that, generically, one can recover $A$ from $\mathrm{Sym}^{2}(A)$ up to a non-zero scalar, which is precisely the level of ambiguity relevant for projective representations.
Concretely, for $q=7$ the script runs a small number of random trials. For each trial it chooses a random $A\in\mathrm{GL}_{2}(7)$, computes $M=\mathrm{Sym}^{2}(A)$, and then searches for all $A^{\prime}\in\mathrm{GL}_{2}(7)$ with $\mathrm{Sym}^{2}(A^{\prime})=M$. Among the candidates it checks whether there is an $A^{\prime}$ which is a scalar multiple of $A$. In all trials this is the case. For example, in one run the script reports
$A=\begin{pmatrix}6&2\\ 2&4\end{pmatrix},\qquad A^{\prime}=\begin{pmatrix}1&5\\ 5&3\end{pmatrix}$,
with the property that $\mathrm{Sym}^{2}(A^{\prime})=\mathrm{Sym}^{2}(A)$. A quick check in $\mathbb{F}_{7}$ shows that $A^{\prime}=\lambda A$ with $\lambda=6$, since
$6\cdot A^{\prime}=6\begin{pmatrix}1&5\\[1.0pt] 5&3\end{pmatrix}=\begin{pmatrix}6&30\\ 30&18\end{pmatrix}\equiv\begin{pmatrix}6&2\\ 2&4\end{pmatrix}\pmod{7}$.
In other trials the script similarly finds that $A$ is determined up to a non-zero scalar by the matrix $\mathrm{Sym}^{2}(A)$.
This toy example thus provides an explicit, concrete illustration of the reconstruction step used in the proof of Theorem 5.1. In that theorem, $W$ is a more complicated tensor product of polynomial modules and the extraction of $A(g)$ from $M_{W}(g)$ uses the known Schur functor structure of each factor $L(\lambda^{(t)})$, rather than brute force. Nevertheless, the small-scale experiment for $\mathrm{Sym}^{2}(V)$ in dimension $2$ confirms the underlying principle: the entries of $M_{W}(g)$ are polynomial expressions in the entries of $A(g)$, and in multiplicity-free situations this information is sufficient to recover $A(g)$ up to scalars (and, in the general setting of the paper, up to field automorphisms).
7.5. Discussion
Although Theorems 4.2, 4.4 and 5.1 are proved purely theoretically, the experiments above provide a useful independent validation of the eigenvalue labelling mechanism and of the reconstruction step in a few small but representative cases. They show that:
  • the base-$q$ injectivity lemma (Lemma 3.1) behaves as predicted for various small triples $(q,d,C)$ with C<q1C<q-1;
  • the exponent formula $E(\nu)=\sum_{i}c_{i}(\nu)q^{i-1}$ correctly encodes the eigenvalues of a Singer cycle on symmetric powers of the natural module, and the digits recovered from discrete logarithms match the expected weight patterns;
  • the labeling strategy scales efficiently to large groups such as $SL_{10}(2^{16})$, identifying weights in millisecond time without the need for explicit field arithmetic in $\mathbb{F}_{2^{160}}$, as demonstrated in Section 7.2;
  • in multiplicity-free modules such as $\mathrm{Sym}^{k}(V)$ for $k\leq 3$ and $(q,d)=(7,3)$, the spectrum of a Singer cycle is indeed simple, and the map from weight patterns to eigenvalues is a bijection;
  • in the toy example $d=2$ with $W=\mathrm{Sym}^{2}(V)$ over $\mathbb{F}_{7}$, the matrix $\mathrm{Sym}^{2}(A)$ determines $A$ up to a non-zero scalar for random choices of $A\in\mathrm{GL}_{2}(7)$, in line with the reconstruction philosophy of Step 4 in the proof of Theorem 5.1.
From the algorithmic point of view, these examples illustrate on a small scale how the rewriting algorithm of Theorem 5.1 uses the eigenvalues of a Singer cycle to label basis vectors by weight patterns and thereby recover the underlying natural module, and how polynomial relations between the entries of $M_{W}(g)$ and of $A(g)$ can be inverted (at least in low-degree cases) to reconstruct $A(g)$ up to scalars. The experiments are intentionally limited to small parameters and to symmetric powers, where explicit matrix computations are still feasible in SageMath. In future work, it would be interesting to complement these tests with analogous computations for more complicated Schur functors and tensor products, and for modules which are not multiplicity-free, in order to explore empirically how far the simple-spectrum and reconstruction properties extend beyond the hypotheses of this paper.